# Battery Capacity Knee-Onset Identification and Early Prediction Using Degradation Curvature

### v2

@@ Abstract
Abrupt capacity fade can have a significant impact on performance and safety in battery applications. To address concerns arising from possible knee occurrence, this work aims for a better understanding of their cause by introducing a new definition of capacity knees and their onset. A curvature-based identification of a knee and its onset is proposed, which relies on a discovered distinctly fluctuating behavior in the transition between an initial and a final stable acceleration of the degradation. The method is validated on both synthetic degradation data and experimental degradation data of two different battery chemistries and is also benchmarked to the state-of-the-art knee identification method in the literature. The results demonstrate that our proposed method could successfully identify capacity knees when the state-of-the-art knee identification method failed. Furthermore, a significantly strong correlation is found between knee and end of life (EoL) and almost equally strong between knee onset and EoL. The results can contribute to a better understanding of capacity knees and the proposed capacity knee identification method can be potentially used for systematically evaluating the knee prediction performance of both model-based methods, and data-driven methods and facilitate better classification of retired automotive batteries from safety and performance perspectives.
@@
A curvature approximation method to measure the rate of change of capacity fade rate is proposed.
A new oscillatory degradation phenomenon is discovered after applying the proposed curvature approximation method.
The new oscillatory degradation phenomenon is used to identify knee and knee-onset points.
The introduced identification method outperforms the state-of-the-art method on both synthetic and experimental data.
The early knee-onset prediction models are demonstrated to significantly reduce experimental time and cost.
@@ 1 Introduction
Lithium-ion batteries have been widely used as energy storage systems in various applications, such as electric vehicles, and microgrids due to their high power and energy density, rapid response, and long lifetime characteristics schmuch2018performance . However, as a result of a complex interplay of different physical and chemical degradation mechanisms, the performance (e.g., available energy and available power) of lithium-ion batteries gradually degrades over their service lives, where the degradation rate is a nonlinear function of storage and cycling conditions (temperature, state-of-charge (SoC) window, charge/discharge current, energy throughput, etc). In some cases, accelerated capacity fade can not only lead to accelerated performance degradation but also safety issues of a battery martinez2018technical .
@@
In both field data and experimental testing of commercial lithium-ion batteries, it has been observed that the capacity fade exhibits a two-stage behavior, with a slow degradation rate in the first stage and then an accelerated degradation rate in the second stage dubarry2012synthesize he2011prognostics yang2017prognostics . The transition from the first stage to the second stage infers a knee pattern on the capacity fade curve. Generally, the knee occurs within a 70-95% capacity retention window depending on the operating conditions attia2022knees . Furthermore, it has been experimentally demonstrated that reusing lithium-ion batteries in less-demanding second-life applications may not slow down the aging trend once the knee has already occurred martinez2018technical martinez2016evaluation zhang2019accelerated . Therefore, for safety and performance reasons, the occurrence of the knee should be avoided, or at least delayed, to ensure a long battery lifetime, and batteries with knee occurrence should normally be retired immediately from operation, either in their first-life or second-life applications martinez2018technical .
@@
It is common for manufacturers of electric vehicles to provide a battery warranty of 8-10 years, guaranteeing 70-80% of their initial nominal capacity wood2011investigation arrinda2021application . When a battery pack that consists of thousands of battery cells retires from its first life in an electric vehicle, not all battery cells in the battery pack have necessarily reached 70-80% of their initial nominal capacity, and knee at cell level may occur before or after the end of life defined at pack level, due to cell-to-cell variations in the pack baumann2018parameter . Instead of being recycled or disposed of, one desirable option is to repurpose retired batteries to less demanding second-life applications as stationary battery energy storage systems (BESSs) ahmadi2017cascaded . However, to ensure the lifetime is maximized in its second-life applications, a knee occurrence on the capacity fade curve needs to be identified prior to repurposing retired batteries for a second-life application.
@@
To date, several studies have attempted to firstly define the knee and then propose a method to identify it, both in off-line scenarios satopaa2011finding diao2019algorithm fermin2020identification greenbank2021automated , and online scenarios zhang2019accelerated . Diao et al. diao2019algorithm defined the knee as the intersection of two tangent lines at two points (i.e., the points with minimum and maximum absolute slope, respectively) on the capacity fade curve, and then developed an empirical degradation model to characterize the capacity fade curve from experimental data, on which the two points were identified to locate the two tangent lines. However, different degradation models may be required to fit different types of capacity fade curves. Zhang et al. zhang2019accelerated firstly learned a strip-shaped safety zone from experimental data of the height of a peak on the incremental capacity curve, and then the knee was identified as the last cycle of four consecutive cycles beyond the safety zone using quantile regression and Monte Carlo simulation. Although the quantile regression method works with incoming data streams, the identified knees vary with the amount of available data online. In a more recent work by Fermín-Cueto et al. fermin2020identification , the knee was defined as the intersection of two straight lines identified by directly fitting the Bacon Watts model to capacity degradation data. The Bacon-Watts method is simple and robust against noise without superimposing a degradation model, but may not be applicable to all types of capacity fade curves. The bisector method, proposed by Greenbank and Howey greenbank2021automated , first fits the early and late life capacity fade gradients using linear regression. Then the knee is identified as the intersection of an angle bisector of two gradients and the capacity fade curve. However, the bisector method is sensitive to the selection of early and late life capacity fade data, and may therefore not be applicable to all types of capacity fade curves. In summary, the knee identification methods in the aforementioned studies may not always be applicable to different types of capacity fade curves. Moreover, possible physical insights that could be obtained by correlating the occurrence of the knee with changes in underlying degradation mechanisms are not considered in these studies.
@@
    We are the first to propose the use of approximated curvature to measure the rate of change of capacity fade rate in discrete time, and this reveals a new oscillatory degradation phenomenon. This degradation phenomenon occurs during a period that clearly separates the capacity fade curve into three distinct intervals that can be automatically identified offline as well as online with some delays. The knee onset corresponds to the start of the oscillatory middle interval and the knee corresponds to the end of this period.
@@
    We validate the effectiveness of our proposed method on two experimental datasets of two different cathode chemistries (i.e., LFP and NMC) and one synthetic dataset, whose knee occurrence is induced by lithium plating, resistance growth, and particle cracking, respectively. Compared to the state-of-the-art method (i.e., the double Bacon-Watts model), our proposed method consistently identifies knee and knee-onset points on capacity fade curves over all three battery datasets while the double Bacon-Watts model completely fails on one battery dataset.
@@
    With strong correlations between knee and knee-onset (identified using our proposed identification method) and the end of life, the knee-onset can provide a much earlier warning of accelerated degradation. Therefore, we demonstrate that learning early-prediction models for battery knee-onset prediction significantly reduces experimental time and its associated cost. Moreover, knee-onset prediction can potentially benefit several industrial applications with significant economic value (e.g., validation of formation protocols, battery grading, battery replacement planning, and battery repurposing for second-life applications).
@@ 2.1 Double Bacon-Watts model
In order to identify both knee and knee-onset points on capacity fade curves instead of only one knee point, the double Bacon-Watts model is adapted from the Bacon-Watts model fermin2020identification , i.e.,
@@
where $\alpha_{0}$ is an intercept; $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ denote the slopes of the intersecting lines; $\gamma$ determines the abruptness of the transition and is set to a low value to achieve an abrupt transition; $Z$ is a zero mean, normally distributed random variable; $x_{0}$ and $x_{2}$ are two transition points, i.e., knee and knee-onset.
@@
$\alpha_{1}$ Slope of the intersecting line 1 -1e-4 Ah/cycle
$\alpha_{2}$ Slope of the intersecting line 2 -1e-4 Ah/cycle
@@ 2.2 Matrix profile
Here, we first introduce all the necessary definitions related to matrix profile yeh2016matrix , which will be used later in our proposed method:
@@ 2.3 Scalable time series anytime matrix profile
To efficiently compute these two time series, i.e., the matrix profile and matrix profile index, the algorithm that we adopt in this work is called Scalable Time series Anytime Matrix Profile (STAMP) yeh2016matrix . STAMP is a time-series all-pairs one-nearest-neighbor search algorithm that uses the Fast Fourier Transform for speed and scalability. There are only two input parameters, i.e., the time series $\bm{Y}$, and a subsequence length $L_{1}$, where $L_{1}$ is the desired length of the time series pattern to search for. A pseudocode representation of the STAMP is given in Algorithm 1.
Algorithm 1 STAMP
1:Input: 1. $\bm{Y}$: time series with the length of $M$
2:Input: 2. $L_{1}$: length of the subsequence
3:Initialize an empty matrix profile $\bm{P}_{\bm{ZZ}}$, an empty matrix profile index $\bm{I}_{\bm{ZZ}}$, and an index array $\bm{I}_{idx}=[1:M-L_{1}+1]$
4:for $i_{idx}$ in $\bm{I}_{idx}$ do
5:     Calculate distance profiles $\bm{D}=\textbf{MASS}(\bm{Z}(i_{idx}),\bm{Y})$ FastestSimilaritySearch
6:     Calculate pairwise minimum for each element in $\bm{D}$, $\bm{P}_{\bm{ZZ}}$, $\,\bm{I}_{\bm{ZZ}}$ = $\textbf{ElementWiseMin}(\bm{P}_{\bm{ZZ}},\bm{I}_{\bm{ZZ}},\bm{D},i_{idx})$
7:end for
8:Output: matrix profile $\bm{P}_{\bm{ZZ}}$ and its associated matrix profile index $\bm{I}_{\bm{ZZ}}$.
@@ 2.4 Arc curve
Here we introduce all the necessary definitions related to the arc curve, which represents the likelihood of a regime change at each location gharghabi2017matrix .
@@ 2.5 Fast low-cost unipotent semantic segmentation
To produce the third time series, i.e., the Corrected Arc Curve (CAC), we adopt the Fast Low-cost Unipotent Semantic Segmentation (FLUSS) gharghabi2017matrix . The FLUSS has only two input parameters, i.e., the matrix profile index $\bm{I}_{\bm{ZZ}}$ and the subsequence length $L_{2}$.
@@
10:IAC=parabolic curve of length $\mathrm{IAC}=\text{parabolic curve of length }n\text{ and height }n/2$ gharghabi2017matrix
@@ 2.6 Regime extracting algorithm
With the advantage of having only two input parameters, i.e., the subsequence length $L_{2}$, and the number of states $N_{S}$, we adopt the regime extracting algorithm (REA) gharghabi2017matrix to extract the locations of the state changes from the CAC, i.e., $\bm{B}$.
@@ 3 Methods
Researchers from different areas come across knee identification problems, in which knees can be detected in either an ad-hoc manner or with a general tool satopaa2011finding . The concept of a knee in the battery field generally relates to the degradation rate on the capacity fade curve. The degradation rate in terms of capacity fade is the result of the convolution of various underlying degradation mechanisms and possible interactions between them. Extrinsic factors, such as the sequence of aging tests, may influence the degradation rate in terms of capacity fade, especially at high C-rates and long duration of continuous cycling raj2020investigation . Intrinsic factors, such as battery chemistry, and manufacturing variances, also have an impact on the degradation rate bach2016nonlinear . To design a general knee identification method, a consistent knee definition that is applicable to batteries of any chemistry and a wide range of operating conditions is required.
To measure the rate of change of the capacity fade, we first introduce the concept of curvature, which is a mathematical measure of the amount by which a curve deviates from being a straight line satopaa2011finding . For a continuous function $f$, the curvature $\kappa(x)$ of $f$ at any point, is defined as
@@
Although a knee can be mathematically well-defined as the point of maximum curvature for continuous functions satopaa2011finding , it is in practice challenging to accurately identify the knee using Eqn (2) as the capacity fade data is sampled and noisy. Therefore, the curvature needs to be approximated before the knee identification on discrete data.
@@ 3.1 Degradation curvature approximation
Given a sequence of measured capacity data from a lithium-ion cell, $\{x_{i},y_{i}\}_{i=1}^{N}$, where $x_{i}\in\mathbb{R}_{0}^{+}$ is the number of cycles that the battery has been used, $y_{i}\in\mathbb{R}^{+}$ is the discharge capacity measured per cycle, and $N$ is the number of sampled capacity points in the set. As an example, a set of discrete capacity fade data points of a sample cell, [b1c0] in the Toyota Research Institute dataset severson2019data , is illustrated in Fig. 1a.
@@
The lithium-ion cell has knee occurrence on its capacity fade curve before the end of the experiment.
@@
    As a low-pass filter that utilizes the local least-square polynomial approximation, the Savitzky-Golay filter provides competitive denoising performance savitzky1964smoothing . Additionally, it is more computationally efficient than other smoothing techniques with the potential for real-time applications. Therefore it is used to smooth the normalized data points $\{x_{i},y_{{n}_{i}}\}_{i=1}^{N}$. The resulting set of smoothed data points $\{x_{i},y_{{sn}_{i}}\}_{i=1}^{N}$ is then used in the next step.
@@
The resulting approximated curvature of the example cell in Fig. 1a is illustrated in Fig. 1b. Contrary to what one would perhaps expect from Fig. 1a, having most fluctuations at the beginning, we observe a very stable start of the curvature followed by a distinct period of large fluctuations. We interpret this as the battery degradation going through three consecutive aging states. First, we have a stable state, where we can expect solid electrolyte interphase (SEI) to play a central role in aging. Then comes a transition phase where a new process (i.e., loss of delithiated negative electrode active material) has started to be significant and causing fluctuations in the curvature, followed by a more stable degradation state, where another new process dominates, i.e., lithium plating.
@@
Figure 1: The curvature approximation results of a sample cell [b1c0] in the Toyota Research Institute dataset severson2019data .
As a matter of fact, drastic changes of system states from one stable state to another stable state through a critical transition state can also be observed in other complex dynamical systems in ecology, biology, economics, and other fields scheffer2009early liu2015identifying . Since we know that the fade curve exhibits accelerated aging in the last phase and a stable degradation (after the low-pass filtering) in the first phase, it seems rational to regard the initiation of the transition phase as the knee onset and then define the end of this phase as the knee point.

### v3

@@ Abstract
Abrupt capacity fade can have a significant impact on performance and safety in battery applications. To address concerns arising from possible knee occurrence, this work aims for a better understanding of their cause by introducing a new definition of capacity knees and their onset. A curvature-based identification of a knee and its onset is proposed, which relies on the discovery of a distinctly fluctuating behavior in the transition between an initial and a final stable acceleration of the degradation. The method is validated on experimental degradation data of two different battery chemistries, synthetic degradation data, and is also benchmarked to the state-of-the-art knee identification method in the literature. The results demonstrate that our proposed method could successfully identify capacity knees when the state-of-the-art knee identification method failed. Furthermore, a significantly strong correlation is found between knee and end of life (EoL) and almost equally strong between knee onset and EoL. As the method does not require the full capacity fade curve, this opens up online knee-onset identification as well as knee and EoL prediction.
@@
A curvature approximation method to measure the rate of capacity fade is proposed.
A new oscillatory degradation phenomenon is discovered.
The oscillatory degradation phenomenon is used to identify knee and knee-onset.
The method outperforms the state-of-the-art method on both synthetic and real data.
The method can be used for online knee-onset identification and knee prediction.
@@ 1 Introduction
Lithium-ion batteries have been widely used as energy storage systems in various applications, such as electric vehicles and microgrids, due to their high power and energy density, rapid response, and long lifetime characteristics [1]. However, as a result of a complex interplay of different physical and chemical degradation mechanisms, the performance (e.g., available energy and available power) of lithium-ion batteries gradually degrades over their service lives, where the degradation rate is a nonlinear function of storage and cycling conditions (temperature, state-of-charge (SoC) window, charge/discharge current, energy throughput, etc). In some cases, accelerated capacity fade can not only lead to accelerated performance degradation but also battery safety issues [2].
@@
In both field data and experimental testing of commercial lithium-ion batteries, it has been observed that the capacity fade often exhibits a two-stage behavior, with a slow degradation rate in the first stage and then an accelerated degradation rate in the second stage [3] [4] [5]. The transition from the first stage to the second stage infers a knee appearance on the capacity fade curve. Generally, the knee occurs within a 70-95% capacity retention window depending on the operating conditions [6]. Furthermore, it has been experimentally demonstrated that reusing lithium-ion batteries in less-demanding second-life applications may not slow down the aging trend once the knee has already occurred [2] [7] [8]. Therefore, for safety and performance reasons, the occurrence of the knee should be avoided, or at least delayed, to ensure a long battery lifetime. Batteries with knee occurrence should normally be retired immediately from operation, no matter if it is in their first-life or later-life application [2].
@@
It is common for manufacturers of electric vehicles to provide a battery warranty of 8-10 years, guaranteeing 70-80% of their initial nominal capacity [9] [10]. When a battery pack that consists of thousands of battery cells retires from its first life in an electric vehicle, not all battery cells in the battery pack have necessarily reached 70-80% of their initial nominal capacity, and knee at cell level may occur before or after the end of life defined at pack level, due to cell-to-cell variations in the pack [11]. Instead of being recycled or disposed of, one desirable option is to repurpose retired batteries to less demanding second-life applications as stationary battery energy storage systems (BESSs) [12]. However, to ensure the lifetime is maximized in its second-life application, a knee occurrence on the capacity fade curve needs to be identified prior to repurposing the retired batteries for a second-life application.
@@
To date, several studies have attempted to firstly define the knee and then propose a method to identify it, both in off-line scenarios [13] [14] [15], and online scenarios [8] [16] [17]. Diao et al. [13] defined the knee as the intersection of two tangent lines at two points (i.e., the points with minimum and maximum absolute slope, respectively) on the capacity fade curve, and then developed an empirical degradation model to characterize the capacity fade curve from experimental data, on which the two points were identified to locate the two tangent lines. However, different degradation models may be required to fit different types of capacity fade curves. Fermín-Cueto et al. [14] also defined the knee as the intersection of two straight lines identified by directly fitting the Bacon Watts model to capacity degradation data. The Bacon-Watts method is simple and robust against noise without superimposing a degradation model, but may not be applicable to all types of capacity fade curves. The bisector method, proposed by Greenbank and Howey [15], first fits the early and late life capacity fade gradients using linear regression. Then the knee is identified as the intersection of an angle bisector of two gradients and the capacity fade curve. However, the bisector method is sensitive to the selection of early and late life capacity fade data, and may therefore not be applicable to all types of capacity fade curves. Zhang et al. [8] firstly learned a strip-shaped safety zone from experimental data of the height of a peak on the incremental capacity curve, and then the knee was identified as the last cycle of four consecutive cycles beyond the safety zone using quantile regression and Monte Carlo simulation. Although the quantile regression method works with incoming data streams, the identified knees vary with the amount of available online data. Sohn et al. [16] proposed a convolutional neural networks model to extract temporal features from time-series data to predict the number of cycles left to the knee point. However, their method requires extensive knee labeling beforehand for the model training process using the Bacon-Watts model. In a very recent work by Costa et al. [17], a transformer-based deep learning model integrated with incremental capacity analysis was proposed for knee identification and degradation diagnosis. The knee identification was formulated into a binary classification problem, i.e., the model predicts whether or not a knee will occur within a window size of 800 cycles. Although the model can quantify degradation modes in addition to the knee indication within 800 cycles, it also requires knee labeling beforehand and does not identify the exact knee point on the capacity fade curve, which is often preferable in practical applications, such as in the classification of retired electric vehicle batteries. To summarize, the aforementioned knee identification methods can be divided into two categories. One is based on finding the intersection between a straight line approximating the early fade and a line following the rate of fade after a knee [13] [14] [15]. The disadvantages of this approach are that, depending on the shape of the capacity fade curve, they may fail, and they cannot be used for online identification and prediction as they need more or less the complete fade curve. The other category is methods based on machine learning models [8] [16] [17]. These methods can be used for online identification and prediction, but they require large amounts of labeled data and other inputs than only capacity.
@@
    We are the first to propose the use of approximated curvature to measure the rate of change of capacity fade rate in discrete time, and this reveals a new oscillatory degradation phenomenon. This degradation phenomenon occurs during a period that clearly separates the capacity fade curve into three distinct intervals that can be automatically identified offline as well as online with some delay. The knee onset corresponds to the start of the oscillatory middle interval and the actual knee corresponds to the end of this period.
@@
    We validate the effectiveness of our proposed method on two experimental datasets of two different cathode chemistries (i.e., LFP and NMC) and one synthetic dataset, the three datasets covering knee occurrence induced by lithium plating, resistance growth, and particle cracking. Compared to the state-of-the-art method (i.e., the double Bacon-Watts model), our proposed method consistently identifies knee and knee-onset points on capacity fade curves over all three battery datasets while the double Bacon-Watts model completely fails on one of the battery datasets.
@@
    With strong correlations between knee and knee-onset (identified using our proposed identification method) and the end of life, the knee-onset can provide an early warning of accelerated degradation. Therefore, we demonstrate that learning early-prediction models for battery knee-onset prediction can significantly reduce experimental time and its associated costs. Moreover, knee-onset prediction can potentially benefit several industrial applications with significant economic value (e.g., validation of formation protocols, battery grading, battery replacement planning, and battery repurposing for second-life applications).
@@ 2.1 Double Bacon-Watts model
In order to identify both knee and knee-onset points on capacity fade curves instead of only one knee point, the double Bacon-Watts model is adapted from the Bacon-Watts model [14], i.e.,
@@
where $\alpha_{0}$ is an intercept; $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ denote the slopes of the intersecting lines; $\gamma$ determines the abruptness of the transition and is set to a low value (i.e., $10^{-8}$) to achieve an abrupt transition; $Z$ is a zero mean, normally distributed random variable; $x_{0}$ and $x_{2}$ are two transition points, i.e., knee-onset and knee.
@@
$\alpha_{1}$ Slope of the intersecting line 1 $-10^{-4}$ Ah/cycle
$\alpha_{2}$ Slope of the intersecting line 2 $-10^{-4}$ Ah/cycle
@@ 2.2 Matrix profile
Here, we first introduce all the necessary definitions related to matrix profile [18], which will be used later in our proposed method:
@@ 2.3 Scalable time series anytime matrix profile
To efficiently compute these two time series, i.e., the matrix profile $\bm{P}_{\bm{ZZ}}$ and matrix profile index $\bm{I}_{\bm{ZZ}}$, the algorithm that we adopt in this work is called Scalable Time series Anytime Matrix Profile (STAMP) [18]. STAMP is a time-series all-pairs one-nearest-neighbor search algorithm that uses the Fast Fourier Transform for speed and scalability. There are only two input parameters, i.e., the time series $\bm{Y}$, and a subsequence length $L_{1}$, where $L_{1}$ is the desired length of the time series pattern to search for.
@@ 2.4 Arc curve
Here we introduce all the necessary definitions related to the arc curve, which represents the likelihood of a regime change at each location [19].
@@ 2.5 Fast low-cost unipotent semantic segmentation
To produce the third time series, i.e., the Corrected Arc Curve (CAC), we adopt the Fast Low-cost Unipotent Semantic Segmentation (FLUSS) algorithm [19]. FLUSS (see Algorithm 1) has only two input parameters, i.e., the matrix profile index $\bm{I}_{\bm{ZZ}}$ and the subsequence length $L_{2}$.
@@
10:IAC=parabolic curve of length $\mathrm{IAC}=\text{parabolic curve of length }n\text{ and height }n/2$ [19]
@@ 2.6 Regime extracting algorithm
With the advantage of having only two input parameters, i.e., the subsequence length $L_{2}$, and the number of states $N_{S}$, we adopt the regime extracting algorithm (REA) [19] to extract the locations of the state changes from the CAC, i.e., $\bm{B}$.
@@ 3 Methods
Researchers from different areas come across knee identification problems, in which knees can be detected in either an ad-hoc manner or with a general tool [20]. The concept of a knee in the battery field generally relates to the degradation rate on the capacity fade curve. The degradation rate in terms of capacity fade is the result of the convolution of various underlying degradation mechanisms and possible interactions between them. Extrinsic factors, such as the sequence of aging tests, may influence the degradation rate in terms of capacity fade, especially at high C-rates and long duration of continuous cycling [21]. Intrinsic factors, such as battery chemistry, and manufacturing variances, also have an impact on the degradation rate [22]. To design a general knee identification method, a consistent knee definition that is applicable to batteries of any chemistry and a wide range of operating conditions is required.
To measure the rate of change of the capacity fade, we first introduce the concept of curvature, which is a mathematical measure of the amount by which a curve deviates from being a straight line [20]. For a continuous function $f$, the curvature $\kappa(x)$ of $f$ at any point, is defined as
@@
Although a knee can be mathematically well-defined as the point of maximum curvature for continuous functions [20], it is in practice challenging to accurately identify the knee using Eqn (2) as the capacity fade data is sampled and noisy. Therefore, the curvature needs to be approximated before the knee identification on discrete data.
@@ 3.1 Degradation curvature approximation
The starting point is a sequence of measured capacity data from a lithium-ion cell, $\{x_{i},y_{i}\}_{i=1}^{N}$, where $x_{i}\in\mathbb{R}_{0}^{+}$ is the number of cycles that the battery has been used, $y_{i}\in\mathbb{R}^{+}$ is the discharge capacity measured per cycle, and $N$ is the number of sampled capacity points in the set. As an example, a set of discrete capacity fade data points of a sample cell, [b1c0] from the Toyota Research Institute dataset [23], is illustrated in Fig. 1a.
@@
The lithium-ion cell has a knee occurrence on its capacity fade curve before the end of the experiment.
@@
    As a low-pass filter that utilizes the local least-square polynomial approximation, the Savitzky-Golay filter provides competitive denoising performance [24]. Since it is more computationally efficient than many other smoothing techniques with potential for real-time applications, it is used to smooth the normalized data points $\{x_{i},y_{{n}_{i}}\}_{i=1}^{N}$. The resulting set of smoothed data points $\{x_{i},y_{{sn}_{i}}\}_{i=1}^{N}$ is then used in the next step.
@@
The resulting approximated curvature of the example cell in Fig. 1a is illustrated in Fig. 1b. Contrary to what one would perhaps expect from Fig. 1a, having most fluctuations at the beginning, we observe a very stable start of the curvature followed by a distinct period of large fluctuations, eventually followed by an abrupt change to yet another stable period.
@@
Figure 1: The curvature approximation results of a sample cell [b1c0] in the Toyota Research Institute dataset [23].
As a matter of fact, drastic changes of system states from one stable state to another stable state through a critical transition state can also be observed in other complex dynamical systems in ecology, biology, economics, and other fields [25] [26]. Since we know that the fade curve exhibits accelerated aging in the last phase and a stable degradation (after the low-pass filtering) in the first phase, it seems rational to regard the initiation of the transition phase as the knee onset and then define the end of this phase as the knee point.